20 years ago, I was half the age I will be in 10 years. How old am I ?
Answer by jim_thompson5910(35256)

x+10 = 2(x-20)

x + 10 = 2x - 40

10 + 40 = 2x - x

50 = x

x = 50

So you are currently 50 years old. In 10 years you will be 60 years old. 20 years ago, you were 30, which is exactly half of 60.